OpenStudy (anonymous):

Add three and one-eighth plus four and five-eighths. one and one-half seven and three-fourths seven and one-half one and three-fourths

jigglypuff314 (jigglypuff314):

3 + 4 = 7 1/8 + 5/8 = 6/8 and because both 6 and 8 can be divided by 2 6/8 = 3/4